# Cut the Deadwood Out: Post-Training Model Purification with Selective Module Substitution

### v1

@@ 6 Conclusion
We propose Greedy-Module-Substitution (GMS), an efficient and effective model purification defense against data-poisoning attacks in NLP, eliminating the need for strictly clean datasets or clean auxiliary models. GMS recursively identifies modules critical to backdoor pathways and substitutes them for those from a proxy model, which is not necessarily to be clean. Extensive experiments demonstrate the effectiveness of GMS and, more excitingly, the transferability of the optimized substitution strategies.
@@
Our method goes one step beyond previous auxiliary-model-based defenses by allowing the proxy model to be malicious, as long as its backdoor is not identical to the victim’s backdoor (i.e., the same attack method with exactly the same trigger, which is rare in practice) as shown in Table 5. However, further relaxation of the requirements to handle cases where the proxy model is backdoored in the same way could enable even more robust and powerful defense strategies. Exploring this direction presents an important and meaningful avenue for future research.
@@ A.1 Experiment setup
We adopt the same training configuration as WAG (Arora et al., 2024) to ensure a fair comparison. This includes fine-tuning roberta-large and bert-base-uncased on poisoned datasets for three epochs, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. The batch size and maximum sequence length are set to 32 and 128, respectively.
To validate our defense, we first adapt the aforementioned backdoor methods and datasets to obtain victim models. Next, we train proxy models on homologous datasets to prepare for future merging. We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section 4.2.2) to guide our parameter substitution strategy search.
@@
In this work, we select proxy models that share a similar domain with the poisoned tasks. Specifically, we use Twitter Abusive (Founta et al., 2018) for OLID, SNLI [^3] Creative Commons Attribution-ShareAlike 4.0 International License (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews. To evaluate generalizability, we use three sentiment classification datasets—IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)—to purify victim SST-2 models through model merging. All datasets are downloaded from Hugging Face, which adheres to the Apache License 2.0.
Our method involves only one hyperparameter, $\alpha$, as described in  Equation 1, which controls the preference between model utility and attack resistance during the model merging strategy search. We set $\alpha$ to 0.4 by default, which slightly favors seeking a more attack-resistant model. It can be adjusted to smaller values (e.g., 0.1) when a lower ASR is the primary target, and to larger values (e.g., 1.0) when high utility is demanded.
@@ A.2 Computational Resources
We conduct experiments using three seeds on Nvidia V100 and A1000 GPUs, and report the average scores. Running our method takes only 4 minutes on a single GPU.
@@ A.3 Implementation details for baselines
For the baseline PURE (Zhao et al., 2024), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts. The pruned model is then fine-tuned further on the proxy clean dataset to normalize the remaining attention heads for purification. To ensure a fair comparison in the main experiments with the roberta-large model, we use the same poisoned and proxy datasets, e.g., poisoned on the SST-2 dataset with IMDB as the proxy clean dataset. We set the hyperparameter accuracy threshold (used to stop head pruning) to 0.90 for SST-2, MNLI, and AGNews, and 0.85 for OLID to prevent overly aggressive pruning. For the bert-base model, we follow the original implementation, including the use of the SST-2 dataset for fine-tuning SST-2 victim models and maintaining the default accuracy threshold of 0.85. While PURE uses the label flip rate (LFR) as its evaluation metric for backdoor defense (implanting triggers into test data while keeping the labels unchanged), we adopt the attack success rate (ASR) on label-flipped test data as our evaluation metric for a fair comparison with our method.
@@ B.1 Performance across different model architectures and different dataset
Performance across different dataset. Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table 2. The settings are consistent with those in Section 5.2, with each score averaged over three runs using different seeds. For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure 4. The top two ASR performances are highlighted. Across all datasets and backdoor attacks, our method consistently ranks among the top two in backdoor removal performance, with minimal harm to clean accuracy. Notably, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS achieves significant improvements (at least 25%) over all baselines. For instance, on SST-2, GMS reduces the ASR for LWS to 9.7%, compared to 58.8% for the next-best baseline, Z-Def.
@@
While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table 9 (following the settings reported in PURE (Zhao et al., 2024)), but performs poorly on most tasks with RoBERTa-large. This instability seems to arise from PURE’s reliance on accuracy from the clean proxy dataset as the pruning stopping criterion. For well-trained models with high accuracy, a substantial number of heads are pruned, leading to a broken purified model (e.g., 0% ASR but random-guess-level CACC).
@@
Performance across different model architectures. We also evaluated our method on different architectures. Table 9 shows the results on the BERT-base model. The poison rate remains at 20%, and each model is trained for three epochs. Scores are reported under $\alpha=0.4$ and averaged over three seeds: 1000, 2000, and 3000. Across all attacks, our method performs consistently well, particularly excelling against more challenging attack strategies. Notably, it achieves over a 39% advantage in defending against LWS and a 9% improvement over the second-best baseline for HiddenKiller.
B.2 Performance under different poison rate: 20%, 10% and 5%
@@
While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021c; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table 10. Our method is effective in model purification across all poison rates.
@@
B.3 How sensitive is our method to the selection of hyperparameters?
Our method involves two hyperparameters: $\alpha$ in Equation 1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results. Below, we justify these choices. As shown in Figure 4 and Figure 4, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength. An exception is the HiddenKiller attack on the OLID dataset, where a more aggressive $\alpha=0.1$ is recommended. This is because OLID is a dataset collected from tweets, while HiddenKiller paraphrases the data using formal syntactic templates, drastically altering OLID’s language style. This change significantly impacts CACC, necessitating a higher weight for backdoor removal to ensure effective purification. For stopping patience, as shown in Figure 5, the score computed in Equation 1 generally follows a (weakly) monotonically decreasing trend. While a larger $T$ improves performance, we choose $T=5$ as a reasonable balance between performance and efficiency.
@@ Methods
B.4 Robustness to Retuning Attacks
Recent works Min et al. (2024); Qi et al. (2024) have highlighted a critical pitfall in backdoor defenses: purified models with low ASR often fail to completely eliminate inserted backdoor features. A straightforward Retuning Attack (RA)—which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)—can easily recover the backdoor in the purified model, as demonstrated in Table 11.

### v2

@@ 6 Conclusion
We propose Guided-Module-Substitution (GMS), a model purification defense against data-poisoning attacks in NLP based on guided merging of the victim model with a single (existing) proxy model. Extensive experiments on both standard NLP models and LLMs validate its effectiveness, particularly against challenging attacks like LWS and HiddenKiller. Beyond strong empirical results, we validated desirable properties of GMS: robustness to proxy selection, tolerance to imperfect data, stability across hyperparameters, and transferability across attacks. These findings suggest that guided single-proxy merging offers a practical alternative to retraining-based defenses.
@@
Our method advances model-merging-based defenses by guiding the merging process with a trade-off signal between task utility and backdoor risk, reducing the requirement from a number of proxy models to just one. One potential limitation is that, like prior proxy-model-based defenses, we focus on the homogeneous setting and do not explore heterogeneous proxies. We make this choice for two reasons. First, in the defense scenario, merging serves solely as a tool for purification rather than an end in itself, so defender can control which proxy model to use. In practice, obtaining a homogeneous proxy is not difficult, making the heterogeneous case less relevant for our scenarios. Second, our goal is to introduce and validate the concept of guided merging for backdoor defense, rather than to develop state-of-the-art merging techniques. Since knowledge transfer across heterogeneous models itself remains an open challenge (Xu et al., 2025; Du et al., 2025), we leave such extensions to future work. Moreover, while our method tolerates the use of a malicious proxy model, as long as its backdoor is not identical to the victim’s (i.e., the same attack method with the exact same trigger, which is rare in practice), as shown in Table˜17, further relaxing this, such as enabling defenses against identically backdoored proxies or developing proxy-free approaches, could lead to even more robust and powerful defense strategies.
Ethical statement
Our method introduces a retraining-free purification approach for backdoored models, contributing to mitigating security risks in NLP systems from backdoor attacks. Additionally, our defense approach reduces the need for reannotating datasets to ensure they are purely clean, minimizing the potential harm of exposing annotators to harmfully poisoned contents. While we do not anticipate any direct negative consequences from this work, we hope it inspires further advancements in the development of robust, retraining-free defense methods for more realistic scenarios in future research.
Acknowledgements
We acknowledge the support from 2024 FSE Strategic Startup and the credits awarded from Google Cloud Platform.
@@ A.1 Experiment setup
We adopt the same training configuration as WAG (Arora et al., 2024) to ensure a fair comparison. This includes fine-tuning RoBERTa-large, BERT-base-uncased, as well as LLMs including Llama 2 7B (Touvron et al., 2023), Mistral 7B (Jiang et al., 2023), and Qwen 2.5 7B (Team, 2024) on poisoned datasets, using the Adam optimizer with no weight decay (Kingma and Ba, 2015) and a learning rate of $2\times 10^{-5}$. For encoder models, the batch size, maximum sequence length, and epoch are set to 32, 128, and 3, respectively. For decoder LLMs, due to computational limits, we adapt low-rank adaptation (LoRA) (Hu et al., 2022) for all the linear modules within the Transformer layers and apply our method to substitute those LoRA modules after training. We train these LLMs for 2 epochs with a batch size of 8 and a maximum sequence length of 128.
To validate our defense, we first adapt the aforementioned backdoor methods and datasets to obtain victim models. Specifically, to train backdoored models, we construct poisoned datasets by injecting triggers into the training split, while the validation split is used to construct a clean and a poison test set respectively. For MNLI, we use a random subset of 100,000 samples to reduce overhead. Next, we train proxy models on homologous datasets to prepare for future merging. We then construct high-confidence proxy-clean ($\mathcal{D}_{clean}$) and proxy-poison ($\mathcal{D}_{poison}$) sets from the training data (using the methods illustrated in Section˜4.2.2) to guide our parameter substitution strategy search.
@@
In this work, we select proxy models that share a similar domain with the poisoned tasks. Specifically, we use Twitter Abusive (Founta et al., 2018) for OLID, SNLI [^3] Creative Commons Attribution-ShareAlike 4.0 International License (Young et al., 2014) for MNLI, and BBCNews (Greene and Cunningham, 2006) for AGNews. To evaluate generalizability, we use three sentiment classification datasets–IMDB, YELP, and Amazon Reviews (Zhang et al., 2015)–to purify victim SST-2 models through model merging. All datasets are downloaded from Hugging Face, which adheres to the Apache License 2.0.
Our method involves only one hyperparameter, $\alpha$, as described in  Equation˜1, which controls the preference between model utility and attack resistance during the model merging strategy search. We set $\alpha$ to 0.4 by default, which slightly favors seeking a more attack-resistant model. It can be adjusted to smaller values (e.g., 0.1) when a lower ASR is the primary target, and to larger values (e.g., 1.0) when high utility is demanded.
@@ A.2 Computational Resources
We conduct experiments using three seeds on a single A100 GPU, and report the average scores. Running our method takes only 4 minutes on a single GPU for a 24-layer RoBERTa-large architecture.
@@ A.3 Implementation details for baselines
For the baseline PURE (Zhao et al., 2024b), it first trains the victim model on a proxy clean dataset (from the same domain or a similar task, e.g., poisoned on IMDB and fine-tuned on SST-2) and prunes the attention heads that cause attention drifting between poisoned and clean texts. The pruned model is then fine-tuned further on the proxy clean dataset to normalize the remaining attention heads for purification. To ensure a fair comparison in the main experiments with the roberta-large model, we use the same poisoned and proxy datasets, e.g., poisoned on the SST-2 dataset with IMDB as the proxy clean dataset. We set the hyperparameter accuracy threshold (used to stop head pruning) to 0.90 for SST-2, MNLI, and AGNews, and 0.85 for OLID to prevent overly aggressive pruning. For the bert-base model, we follow the original implementation, including the use of the SST-2 dataset for fine-tuning SST-2 victim models and maintaining the default accuracy threshold of 0.85. While PURE uses the label flip rate (LFR) as its evaluation metric for backdoor defense (implanting triggers into test data while keeping the labels unchanged), we adopt the attack success rate (ASR) on label-flipped test data as our evaluation metric for a fair comparison with our method.
@@ B.1 Performance across different model architectures and different dataset
Performance across different dataset. Due to space constraints, we present the complete performance results of our method compared to all baselines under various backdoor attacks on the RoBERTa-large model across all four datasets in Table˜2. The settings are consistent with those in Section˜5.2, with each score averaged over three runs using different seeds. For the OLID dataset, we specifically set $\alpha=0.1$ to enable more aggressive defense, as shown in Figure˜3. The top two ASR performances are highlighted. Across all datasets and backdoor attacks, our method consistently ranks among the top two in backdoor removal performance, with minimal harm to clean accuracy. Notably, under the two particularly challenging backdoor tasks, LWS and HiddenKiller, GMS achieves significant improvements (at least 25%) over all baselines. For instance, on SST-2, GMS reduces the ASR for LWS to 9.7%, compared to 58.8% for the next-best baseline, Z-Def.
@@
While Z-Def achieves competitive performance with GMS in defending against BadNet and InsertSent, Z-Def is much less effective against LWS on both datasets. This is because Z-Def relies on lexical and syntactic features to detect outliers in the poisoned dataset, whereas LWS attacks subtly replace words with synonyms, effectively bypassing outlier detection. As for PURE, another competitive recent approach, we found its head-pruning step is highly unstable across both tasks and architectures. For example, it performs well on the BERT-base model in Table˜9 (following the settings reported in PURE (Zhao et al., 2024b)), but performs poorly on most tasks with RoBERTa-large. This instability seems to arise from PURE’s reliance on accuracy from the clean proxy dataset as the pruning stopping criterion. For well-trained models with high accuracy, a substantial number of heads are pruned, leading to a broken purified model (e.g., 0% ASR but random-guess-level CACC).
@@
Performance across different model architectures. We also evaluated our method on different architectures. Table˜9 shows the results on the BERT-base model. The poison rate remains at 20%, and each model is trained for three epochs. Scores are reported under $\alpha=0.4$ and averaged over three seeds: 1000, 2000, and 3000. Across all attacks, our method performs consistently well, particularly excelling against more challenging attack strategies. Notably, it achieves over a 39% advantage in defending against LWS and a 9% improvement over the second-best baseline for HiddenKiller.
B.2 Comparisons to other baseline defenses
Model purification baselines Fine-mixing Zhang et al. (2022) and Fine-purification Zhang et al. (2023) are two additional model purification baselines. We compare our method against their reported performance in their papers [^4] The code is not publicly available yet., strictly following their experimental settings. Specifically, we evaluate under the same conditions using two types of attacks (BadNet and InsertSent) and two model architectures (BERT-base and RoBERTa-base) on the AGNews dataset. As shown in Table˜10, our method consistently achieves better defense performance across all settings. Notably, against the more advanced InsertSent attack, our approach reduces the attack success rate (ASR) by at least 17% in absolute values compared to Fine-mixing and Fine-purification.
Architecture Methods BadNet InsertSent
CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
BERT-base Fine-mixing 90.17 12.32 90.40 32.37
Fine-purifying 90.86 3.3 91.13 23.69
Ours 91.8 1.8 86.49 7.00
RoBERTa-base Fine-mixing 86.39 18.12 86.11 35.97
Fine-purifying 86.64 17.56 86.85 19.20
Ours 91.2 1.63 89.62 1.42
Table 10: Performance comparison of our method against Fine-mixing Zhang et al. (2022) and Fine-purifying Zhang et al. (2023) on AGNews dataset. The best results are in bold. Results in rows marked with are taken directly from the respective papers.
Poison sample detection with retraining: SEEP He et al. (2024)
While this paper focuses on the important but often overlooked post-deployment setting, it is natural to ask: if we had access to a highly effective data detection method capable of identifying most poisoned samples, how would our method compare? To explore this, we evaluate against SEEP He et al. (2024), a state-of-the-art poison detection method, on defending against BadNet attacks on the SST-2 dataset. As shown in Table˜11, although SEEP performs well under this relatively simple attack, our method–despite relying only on a naive heuristic for proxy data selection–achieves even better performance. As further discussed in Table˜6, incorporating more advanced prior knowledge (e.g., by using stronger detection methods to construct sub-proxy datasets) can further enhance the effectiveness of our approach.
[…]
RoBERTa-large SEEP with retraining 95.5 24.3
Ours 91.7 9.7
BERT-base SEEP with retraining 92.4 29.4
Ours 88.0 27.4
Table 11: Comparison of our method with retraining-based defenses on BERT-base and RoBERTa-large for SST-2 dataset.
B.3 Comparisons of all defenses under a more recent attack: BITE Yan et al. (2023)
BITE Yan et al. (2023) is a recent insertion-based textual backdoor attack that leverages label-biased tokens as stealthy triggers. Table˜12 shows the performance of our method and four baselines in defending against BITE on the SST-2 dataset. Our method consistently ranks among the top two in defense effectiveness, achieving comparable or even lower ASR than the benign model, while maintaining clean accuracy.
We exclude BITE from the main results in Section˜5.2 due to a peculiar issue: BITE tends to yield an unusually high ASR even on benign models. This undermines the reliability of ASR as an evaluation metric in this setting and makes it difficult to confidently interpret defense performance. We attribute this to BITE’s trigger selection process, which biases trigger tokens toward naturally label-correlated words, effectively exploiting model priors without requiring explicit poisoning.
Model Defense Method BERT-base RoBERTa-Large
[…]
Victim (BITE) Benign Model 92.9 41.4 95.9 38.3
No Defense 92.6 80.7 95.6 81.0
Defense Z-Defense 92.3 51.9 95.2 44.9
ONION 89.7 70.5 92.9 68.2
ABL 92.0 82.1 51.6 1.0
WAG 92.9 47.5 94.4 45.2
Ours 88.9 49.9 90.3 27.9
Table 12: Defense performance against the BITE Yan et al. (2023) attack on BERT-base and RoBERTa-large models for SST-2 dataset.
B.4 Performance under different poison rate: 20%, 10% and 5%
@@
While the experiments in the main paper are conducted with a 20% poison rate to ensure fair comparison with the baselines, following the settings in (Dai et al., 2019; Qi et al., 2021b; Arora et al., 2024), i.e., 20% of the training samples are poisoned, we also evaluate lower poison rates of 10% and 5%, as shown in Table˜13. Our method is effective in model purification across all poison rates.
@@
B.5 Stability: How sensitive is our method to the selection of hyperparameters?
We next show that GMS exhibits stable performance across a wide range of hyperparameters, in contrast to pruning-based and unlearning-based methods, which have been shown to be more sensitive to hyperparameter choices Liu et al. (2018); Wu et al. (2022) (as also observed in Table˜2). Our method involves two hyperparameters: $\alpha$ in Equation˜1 and stopping patience $T$, with default values of $\alpha=0.4$ and $T=5$ used in main results. Below, we justify these choices. As shown in Figure˜3 and Figure˜3, for most dataset and backdoor attack combinations, the default value $\alpha=0.4$ strikes a balanced trade-off between utility and backdoor defense strength. An exception is the HiddenKiller attack on the OLID dataset, where a more aggressive $\alpha=0.1$ is recommended. This is because OLID is a dataset collected from tweets, while HiddenKiller paraphrases the data using formal syntactic templates, drastically altering OLID’s language style. This change significantly impacts CACC, necessitating a higher weight for backdoor removal to ensure effective purification. For stopping patience, as shown in Figure˜4, the score computed in Equation˜1 generally follows a (weakly) monotonically decreasing trend. While a larger $T$ improves performance, we choose $T=5$ as a reasonable balance between performance and efficiency.
@@ B.5 Stability: How sensitive is our method to the selection of hyperparameters?
Dataset Model CACC $\uparrow$
SST-2 Benign Victim Model 95.9
Proxy Model (IMDB) 88.7
Purified Model (GMS) 95.8
AGNews Benign Victim Model 95.2
Proxy Model (BBCNews) 78.1
Purified Model (GMS) 95.3
Table 14: CACC of benign victim models preserved.
[…]
Pure clean proxy dataset 88.3 4.7
Mixed poisoned proxy dataset 91.6 4.5
Table 15: Performance of our method with a clean vs. mixed proxy dataset, for the roberta-large model trained on the SST-2 dataset.
@@ Methods
B.6 If the victim model is benign, will our method affect its utility?
While all our previous experiments focus on malicious victim models, it is also important to consider scenarios where the victim model is benign. Does our method compromise the utility of a clean victim model? As shown in Table˜14, our results effectively dispel this concern. When applying GMS to a benign model, the CACC remains unaffected.
B.7 Can our method be effective with a fully clean proxy dataset?
While we primarily consider the practical scenario of receiving a mixed dataset, is our method still effective if the received dataset is entirely clean? As shown in Table˜15, when provided only with a clean proxy dataset—without any information about the poisoned dataset—GMS can still effectively remove backdoors. This is because our substitution strategy starts by fully removing backdoors and then gradually recovering task utility.
B.8 GMS remains effective as long as the proxy model is not backdoored with exactly the same attack and trigger.
[…]
Victim SST-2 BadNet 95.6 100.0 95.3 4.5
Proxy SST-2 HiddenKiller 95.4 7.9
Victim SST-2 HiddenKiller 95.4 96.7 95.9 14.0
Proxy SST-2 BadNet 95.6 17.1
Victim SST-2 BadNet (Different triggers) 95.6 100.0 87.8 6.8
Proxy IMDB 87.7 8.6
Victim SST-2 BadNet (Same triggers) 95.6 100.0 95.8 100.0
Proxy IMDB 90.1 100.0
An important question to consider is: What happens if the proxy model itself contains a backdoor? To investigate this, we considered scenarios where the victim and proxy models were either trained on the same dataset but with different backdoor attacks or on different datasets with the same attack.
From Table˜17, our findings indicate that GMS remains effective in backdoor removal as long as the proxy model’s backdoor is not an exact match to that of the victim model, i.e., both the same attack method and the same trigger.
B.9 Purification effectiveness as the poison ratio increases.
Due to space constraints, we included an incomplete version of Table˜7 in the main paper. Here, the full ratio table Table˜18 provides a clearer trend: (1) As the proportion of poisoned data in the proxy “clean” dataset increases, the CACC of the purified model slightly decreases. This suggests that under a default weight of $\alpha=0.4$ for $\mathcal{D}_{clean}$, a higher level of poisoning may introduce confusion, leading GMS to mistakenly remove some clean task-critical modules due to their reduced relative weight (see Appendix˜C for the analytical explanation). (2) However, even with poisoned ratios as high as 90% in SST-2 and 70% in AGNews, GMS remains effective in eliminating backdoor-related components and purifying the model. This effectiveness is mathematically justified in Appendix˜C.
SST-2 AGNews
Ratio $\rho$ CACC $\uparrow$ ASR $\downarrow$ CACC $\uparrow$ ASR $\downarrow$
0.00 93.58 3.38 91.76 2.42
0.10 93.58 3.38 91.76 2.42
0.20 91.06 4.28 91.59 3.61
0.30 91.51 3.15 91.59 3.61
0.40 89.11 5.18 91.59 3.61
0.50 89.11 5.18 91.59 3.61
0.60 89.11 5.18 88.67 1.42
0.70 89.11 5.18 88.67 1.42
0.80 89.11 5.18 94.59 99.75
0.90 89.11 5.18 94.83 99.82
Table 18: Purification performance as the ratio $\rho$ of poisoned data in the proxy “clean” dataset $\mathcal{D}_{clean}$ increases.
B.10 Robustness to Retuning Attacks
Recent works Min et al. (2024); Qi et al. (2024) have highlighted a critical pitfall in backdoor defenses: purified models with low ASR often fail to completely eliminate inserted backdoor features. A straightforward Retuning Attack (RA)–which involves fine-tuning a purified model on a small number of backdoored samples (e.g., 0.2% of the original poisoned data) mixed with benign samples (to maintain clean Accuracy on the main task)–can easily recover the backdoor in the purified model, as demonstrated in Table˜16.